# Ramification filtration via deformations, II

### v3

@@ 3.4. Relation between $\mathcal{L}^{(v_{0})}$ and $\mathcal{L}^{[v_{0}]}$
where $\iota\in\mathfrak{A}^{0}$, $-\iota\geqslant-qp^{M-1}(|\bar{a}_{1}|+|\bar{a}_{2}|)+p^{M-1-A}b^{*}$. (Recall that $U(1,A)=p^{-A}$, cf. Section 1.)
Note that if the term $Q$ from (3.6) is such that
$|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant(b^{*}/q)U(w,A)$
then $b^{*}/q$ could be replaced by $v_{0}$, cf. Section 2.1, and therefore
$p^{A}\left[[D_{\bar{a}_{1}\bar{n}_{1}}],[D_{\bar{a}_{2}\bar{n}_{2}}]\right]\in\mathcal{L}_{\mathcal{N}}(w+1)\subset\sum_{w_{1}+w_{2}=w+1}[\mathcal{L}_{\mathcal{N}}(w_{1}),\mathcal{L}_{\mathcal{N}}(w_{2})]\,.$
By inductive assumption $b_{w})$ this term belongs to the $W_{M}(k)$-extension of scalars of
@@
This implies that the coefficients for $p^{A}t_{1}^{-\iota}$ in the non-linear part of $y$ with $-\iota\leqslant-p^{M-1}b^{*}U(w,A)$ belong to $[\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w+1}(\mathcal{L})$ and could be ignored. Indeed, since $-\iota\geqslant-qp^{M-1}(|\bar{a}_{1}|+|\bar{a}_{2}|+p^{M-1}b^{*}p^{M-1-A}$ it holds that $|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant(b^{*}/q)(U(w,m)+b^{*}p^{-m})$. All other terms from the non-linear part will contribute just to smaller upper numbers and also can be ignored. As a result, it remains to care just about the linear part.
If $y_{\iota}=p^{m}t_{1}^{-\iota}l_{\iota}$ is a term from the linear part $L(y)$ such that $\iota>0$ then $\iota\in\mathfrak{A}^{+}(p)$. Therefore, such terms do not appear with $t_{1}^{-\iota p^{r}}$, $r\geqslant 1$. By Prposition 2.4 for different $\iota$, these terms give independent contribution to the upper ramification number in (3.4).
Consider the term $p^{m}t^{-\iota}D_{a0}$ with $-\iota=-aqp^{M-1}+b^{*}p^{M-1}U(w,m)$. If it satisfies the condition $(\bullet_{1})$ then $\iota>0$ and it contributes to the upper number $qp^{i}a-b^{*}(p^{i}U(w,m))$ where $p^{i}U(w,m)\in\mathbb{Z}^{+}(p)$ with $i\geqslant 0$.
Therefore, $qp^{i}a-b^{*}p^{i}U(w,m)>0$ and by Sec.2.1, it is
$\geqslant qp^{i}v_{0}U(w,m)-b^{*}p^{i}U(w,m)=(qv_{0}-b^{*})p^{i}U(w,m)\geqslant qv_{0}-b^{*}\,.$
As a result, $p^{m}D_{a0}\in\mathcal{L}^{(v_{0})}+C_{w+1}(\mathcal{L})$. Part $b_{w+1}$) is proved.
@@
Applying $\bar{\mathcal{V}}^{0}$ to (3.2) we obtain the following congruence in $\bar{\mathcal{N}}^{{\dagger}}[0]$ modulo extension of scalars of the ideal [¯$[\bar{\mathcal{L}}^{{\dagger}[v_{0}]},\bar{\mathcal{L}}^{{\dagger}}]+C_{w+1}(\bar{\mathcal{L}}^{{\dagger}})$:
@@
$\Omega\left(\sigma^{M-1}\bar{e}_{1}^{(q)}\right)\equiv\sigma(\bar{X})\circ\sigma^{M-1}\bar{e}_{1}^{(q)}\circ(-\bar{X})\,\mathrm{mod}\,\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{sep}\,.$
@@
Choose the $\iota$-components $\bar{Y}_{\iota}$ such that $\sigma\bar{Y}_{\iota}-\bar{Y}_{\iota}=t_{1}^{-\iota}(A^{{\dagger}}-\mathrm{id})(D_{\iota 0})$ and $\bar{Y}=\sum_{\iota}\bar{Y}_{\iota}$. All $\iota$ give independent contributions to upper ramification numbers. Take $\iota\in\mathfrak{A}_{1}^{+}(p)$, then $\sigma(\bar{Y}_{\iota})-\bar{Y}_{\iota}=t_{1}^{-\iota}B^{{\dagger}}(D_{\iota 0})$.
If $\iota=p^{m}(qp^{M-1}\alpha-p^{M-1-A}b^{*})$, cf. Prop. 2.4b), then it gives the upper ramification $q(p^{A}\alpha)-b^{*}$. On the other hand, $\alpha>p^{-A}b^{*}/q$ is equivalent to $\alpha\geqslant p^{-A}v_{0}$ or $qp^{A}\alpha-b^{*}\geqslant qv_{0}-b^{*}$. So, $\bar{\mathcal{V}}B^{{\dagger}}(D_{\iota 0})\in\bar{\mathcal{L}}^{(v_{0})}$.
@@ 3.6. Interpretation in terms of deformations
We start, as earlier, with Lie algebra $\mathcal{L}$ and $e=e(t)\in\mathcal{L}_{O_{M}(\mathcal{K})}$ such that (for simplicity) $e=\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}$. Choose $f\in\mathcal{L}_{sep}$ such that $\sigma f=e\circ f$ and fix identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$ via $\tau\mapsto(-f)\circ\tau(f)$, $\tau\in\mathcal{G}_{<p,M}$. As earlier, define the filtration $\mathcal{L}(p)$ and set ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$. Then ¯$\bar{\mathcal{L}}_{W_{M}(k)}$ is generated by $D_{an}$, where $a\in\mathbb{Z}^{0}(p)\cap[0,(p-1)v_{0})$. If $\bar{e}\in\bar{\mathcal{L}}_{O_{M}(\mathcal{K})}$ and $\bar{f}\in\bar{\mathcal{L}}_{sep}$ are the images of $e$ and, resp., $f$ under extensions of scalars of the projection $\mathcal{L}\to\bar{\mathcal{L}}$ then $\tau\mapsto(-\bar{f})\circ\tau(\bar{f})$ determines projection $\bar{\eta}^{0}:\mathcal{G}_{<p,M}\to G(\bar{\mathcal{L}})$. Clearly, $\bar{\eta}^{0}$ is induced by $\eta^{0}$.
As earlier, choose $r^{*}$, $N^{*}$, set $q=p^{N^{*}}$, $b^{*}=r^{*}(q-1)$ and introduce the system of coefficients $\{t^{-\iota}\ |\ \iota\in\mathfrak{A}^{0}\}$ from Section 2.2. Consider the quotient ${\mathcal{L}}^{{\dagger}}$ of $\mathcal{L}$ generated by
$\{D_{an}\ |\ a\in[0,(p-1)v_{0}\}\cup\{D_{\iota n}\ |\ \iota\in\mathfrak{A}^{+}(p)\}\subset\{D_{an}\ |\ a\in\mathbb{Z}^{0}(p)\}\,.$
As earlier, define a filtration on $\mathcal{L}^{{\dagger}}$ and set ¯$\bar{\mathcal{L}}^{{\dagger}}=\mathcal{L}^{{\dagger}}/\mathcal{L}^{{\dagger}}(p)$. Note that by Prop. 2.4, all $D_{\iota n}\in\bar{\mathcal{L}}(p)_{W_{M}(k)}$, and therefore, we have the induced projection $\bar{\mathcal{V}}:\bar{\mathcal{L}}^{{\dagger}}\to\bar{\mathcal{L}}$.
@@
$e^{{\dagger}}=\sum_{a\in[0,(p-1)v_{0})}t^{-qp^{M-1}a}D_{a,M-1}+\sum_{\iota\in\mathfrak{A}^{+}(p)}t^{-\iota}D_{\iota 0}\,.$
Then the image of $\bar{e}^{{\dagger}}$ in ¯$\bar{\mathcal{L}}_{O_{M}(\mathcal{K})}$ coincides with $\sigma^{M-1+N^{*}}(\bar{e})$. Therefore, we can choose $f^{{\dagger}}\in\bar{\mathcal{L}}^{{\dagger}}_{sep}$ such that its image in ¯$\bar{\mathcal{L}}_{sep}$ coincides with $\sigma^{M-1+N^{*}}\bar{f}$. This means that we have epimorphism $\mathcal{G}_{<p,M}\to G(\bar{\mathcal{L}}^{{\dagger}})$ which gives $\bar{\eta}^{0}$ when composed with projection $G(\bar{\mathcal{V}}):G(\bar{\mathcal{L}}^{{\dagger}})\to G(\bar{\mathcal{L}})$.
Now Theorem 3.1 states that a unipotent $\mathbb{Z}/p^{M}$-action determined by $t^{qp^{M-1}}\mapsto t^{qp^{M-1}}\widetilde{\exp}(-t^{b^{*}})^{p^{M-1}}$ induces via formalism of nilpotent Artin-Schreier theory a deformation $B^{{\dagger}}:\mathrm{Ker}\,\bar{\mathcal{V}}\to\mathrm{Ker}\,\bar{\mathcal{V}}$. This deformation allows us to recover a “substantial” part ¯$\bar{\mathcal{L}}^{(v_{0})}$ of the ramification ideal $\mathcal{L}^{(v_{0})}$. It appears as $\bar{\mathcal{V}}(\mathrm{Ker}\,\bar{\mathcal{V}})$, i.e. as a deformation of the “trivial” part $\mathrm{Ker}\,\bar{\mathcal{V}}\subset\mathcal{L}(p)$ via nilpotent operator $B^{{\dagger}}$.
@@ 4.1. Agreements
We need to treat the Campbell-Hausdorff group law via exponentials in the appropriate enveloping algebras $\bar{\mathcal{A}}$ and $\bar{\mathcal{A}}^{{\dagger}}$ of $\bar{\mathcal{L}}$ and, resp., $\bar{\mathcal{L}}^{{\dagger}}$, by involving extension of scalars to $O^{\prime}_{M}(\mathcal{K}^{\prime})$. As a matter of fact, the definition of the action $\Omega$ requires working inside $\widetilde{\mathcal{N}}^{{\dagger}}$ but can be formally treated in $\bar{\mathcal{N}}^{{\dagger}}$, cf. Prop. 2.12. So, it makes sense to work mainly with the Lie algebra ¯$\bar{\mathcal{L}}^{{\dagger}}[0]\simeq\bar{\mathcal{L}}$ and its enveloping algebra $\bar{\mathcal{A}}^{{\dagger}}[0]$.
We are going also to use the advantage of special coefficients from $\mathfrak{A}^{0}$. So, we need just a part of the quotient $(\bar{\mathcal{A}}^{{\dagger}}/\bar{\mathcal{J}}^{{\dagger}p})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ generated by the elements of $\bar{\mathcal{N}}^{{\dagger}}[0]$. (Here $\bar{\mathcal{J}}^{{\dagger}}$ is the augmentation ideal of $\bar{\mathcal{A}}^{{\dagger}}$.) Even more, the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is completely determined by the elements $\bar{\mathcal{V}}^{0}B^{{\dagger}}(D_{\iota 0})$ with $\mathrm{ch}(\iota)=1$. Therefore, it would be reasonable to ignore all $\iota$ with $\mathrm{ch}(\iota)\geqslant 2$, i.e. to work with the part generated by $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}[0]:=\bar{\mathcal{N}}^{{\dagger}}[0,0]\oplus\bar{\mathcal{N}}^{{\dagger}}[1,0]$.
As a result we are going to perform calculations inside the $W_{M}(k)$-module $\bar{\mathcal{A}}^{{\dagger}}_{\leqslant 1}[0]$ generated by all $p^{A}t_{1}^{-\iota}D_{\bar{a}\bar{n}}$, where for $\bar{a}=(a_{1},\dots,a_{r})$ and $\bar{n}=(n_{1},\dots,n_{r})$, $D_{\bar{a}\bar{n}}=D_{a_{1}n_{1}}\ldots D_{a_{r}n_{r}}$, $\iota\in\mathfrak{A}^{0}$, $\mathrm{ch}(\iota)\in\{0,1\}$ and it holds $-\iota\geqslant-qp^{M-1}|\bar{a}|+p^{M-1-A}b^{*}$ (recall that $U(1,A)=p^{-A}$). Since we do not need to follow the effects related to $\mathrm{ch}(\iota)\geqslant 2$, we can agree that $t_{1}^{-\iota}\cdot t_{1}^{-\iota^{\prime}}=0$ if $\iota,\iota^{\prime}\in\mathfrak{A}^{0}_{1}$. We do need to go to $\bar{\mathcal{A}}^{{\dagger}}_{\leqslant 1}[0]$ after applying exponential map to $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}$ but almost immediately we come back to $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}$ and even to $\bar{\mathcal{N}}^{{\dagger}}[1,0]$.

### v5

@@ 3.4. Relation between $\mathcal{L}^{(v_{0})}$ and $\mathcal{L}^{[v_{0}]}$
where $\iota\in\mathfrak{A}^{0}$, $\mathrm{ch}(\iota)=s$ and $-\iota\geqslant-qp^{M-1}(|\bar{a}_{1}|+|\bar{a}_{2}|)+p^{M-1}b^{*}U(s,A)$ with some $s\geqslant 1$. Any such term $Q(\iota,\bar{a}_{1},\bar{a}_{2})$ contributes to the $\iota_{0}$-part iff $\iota=p^{r}\iota_{0}$ and $s=\mathrm{ch}(\iota_{0})=1$.
Lemma 3.6.
a) Any $Q(\iota_{0}p^{u},\bar{a}_{1},\bar{a}_{2})\in\mathcal{L}(w+1)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$.
b) If $Q\in\mathcal{L}_{\mathcal{N}}(w+1)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ then it belongs to $([\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w+1}(\mathcal{L}))_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$.
Proof.
a) The linear $\iota_{0}$-term is
$p^{m}D_{a0}t_{1}^{-qp^{M-1}a+p^{M-1}b^{*}/p^{m}}=p^{m}D_{a0}t_{1}^{-p^{M-m-1}(qp^{m}a-b^{*})}\,.$
It contributes with the upper ramification number $qp^{m}a-b^{*}$ (use the inequality $ap^{m}\geqslant v_{0}$). Consider
$Q(\iota_{0}p^{u},\bar{a}_{1},\bar{a}_{2})=p^{A}t_{1}^{-qp^{M-1}b+p^{M-1-A}b^{*}}[[D_{\bar{a}_{1}\bar{n}_{1}}],[D_{\bar{a}_{2}\bar{n}_{2}}]]$
which contributes to the $\iota_{0}$-part. Here $-b=-p^{m-A}a\geqslant-qp^{M-1}(|\bar{a}_{1}|+|\bar{a}_{2}|)$. Therefore, $|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant p^{m-A}a\geqslant p^{m-A}U(w,m)v_{0}$.
Recall that $U(w,m)=w_{0}/p^{u}+w_{1}/p^{u+1}$, where $u\geqslant 0$, $w=w_{0}+w_{1}$, $m=wu+w_{1}$, $0\leqslant w_{1}<w$.
In the case $m-A\leqslant u$ we have $p^{m-A}U(w,m)=U(w,A)$ and, therefore, $Q(\iota_{0},\bar{a}_{1},\bar{a}_{2})\in\mathcal{L}(w+1)_{O_{M}^{\prime}(\mathcal{K}^{\prime})}$.
If $m-A=u+u_{1}$ with $u_{1}\geqslant 1$ then
$|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant p^{u_{1}}(w_{0}+w_{1}/p)v_{0}>wv_{0}$
and already $[[D_{\bar{a}_{1}\bar{n}_{1}}],[D_{\bar{a}_{2}\bar{n}_{2}}]]\in\mathcal{L}(w+1)_{W(k)}$.
b) We know that $|\bar{a}_{1}|+|\bar{a}_{2}|\geqslant(b^{*}/q)U(w,A)\,$, where $b^{*}/q$ could be replaced by $v_{0}$, cf. Section 2.1, and, therefore, $p^{A}\left[[D_{\bar{a}_{1}\bar{n}_{1}}],[D_{\bar{a}_{2}\bar{n}_{2}}]\right]$ belongs to the $W(k)$-extension of scalars of
$\mathcal{L}_{\mathcal{N}}(w+1)\cap C_{2}(\mathcal{L})\subset\sum_{w_{1}+w_{2}=w+1}[\mathcal{L}_{\mathcal{N}}(w_{1}),\mathcal{L}_{\mathcal{N}}(w_{2})]\,.$
By inductive assumption $b_{w})$ the RHS belongs to
@@
As a result of the above Lemma, the $\iota_{0}$-contribution of the non-linear part of $y$ belongs to $[\mathcal{L}^{(v_{0})},\mathcal{L}]+C_{w+1}(\mathcal{L})$ (and could be ignored modulo $\mathcal{L}^{(v_{0})}_{sep}$).
It remains to note that if $y_{\iota}=p^{m}t_{1}^{-\iota_{0}p^{u}}l_{u}$ is a term from the linear part $L(y)$ then $u=0$. Therefore, $p^{m}D_{a0}$ is the only $\iota_{0}$-contribution from $L(y)$. This proves the property $(\bullet_{1})$ and part $b_{w+1}$) is completely proved.
@@
Applying $\bar{\mathcal{V}}^{0}$ to (3.2) we obtain the following congruence in $\bar{\mathcal{N}}^{{\dagger}}[0]$ modulo the extension of scalars of the ideal [¯$[\bar{\mathcal{L}}^{{\dagger}[v_{0}]},\bar{\mathcal{L}}^{{\dagger}}]+C_{w+1}(\bar{\mathcal{L}}^{{\dagger}})$:
@@
$\Omega\left(\sigma^{M-1}\bar{e}_{1}^{(q)}\right)\equiv\sigma(\bar{X})\circ\sigma^{M-1}\bar{e}_{1}^{(q)}\circ(-\bar{X})\,\mathrm{mod}\,\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{sep}\,$
@@
Choose $\iota$-components $\bar{Y}_{\iota}$ such that $\sigma\bar{Y}_{\iota}-\bar{Y}_{\iota}=t_{1}^{-\iota}\bar{\mathcal{V}}(A^{{\dagger}}-\mathrm{id})(D^{\,{\dagger}}_{\iota 0})$ and $\bar{Y}=\sum_{\iota}\bar{Y}_{\iota}$. As earlier, all $\iota$ give independent contributions to upper ramification.
Take $\iota\in\mathfrak{A}_{1}^{+}(p)$, i.e. $\mathrm{ch}(\iota)=1$, then $\sigma(\bar{Y}_{\iota})-\bar{Y}_{\iota}=t_{1}^{-\iota}B^{\,{\dagger}}(D^{{\dagger}}_{\iota 0})$.
If $\iota=p^{m}(qp^{M-1}\alpha-p^{M-1-A}b^{*})$, cf. Prop. 2.4b), then it gives the upper ramification number $q(p^{A}\alpha)-b^{*}$. On the other hand, $\alpha>p^{-A}b^{*}/q$ is equivalent to $\alpha\geqslant p^{-A}v_{0}$ or $qp^{A}\alpha-b^{*}\geqslant qv_{0}-b^{*}$. So, $\bar{\mathcal{V}}B^{{\dagger}}(D_{\iota 0})\in\bar{\mathcal{L}}^{(v_{0})}$.
@@ 3.6. Interpretation in terms of deformations
We start, as earlier, with the Lie algebra $\mathcal{L}$ and $e=e(t)\in\mathcal{L}_{O_{M}(\mathcal{K})}$ such that (for simplicity) $e=\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}$. Choose $f\in\mathcal{L}_{sep}$ such that $\sigma f=e\circ f$ and fix the identification $\eta^{0}:\mathcal{G}_{<p,M}\simeq G(\mathcal{L})$ via $\tau\mapsto(-f)\circ\tau(f)$, where $\tau\in\mathcal{G}_{<p,M}$. As earlier, define the filtration $\mathcal{L}(w)$ and set ¯$\bar{\mathcal{L}}=\mathcal{L}/\mathcal{L}(p)$. Then ¯$\bar{\mathcal{L}}_{W(k)}$ is generated by $D_{an}$, where $a\in\mathbb{Z}^{0}(p)\cap[0,(p-1)v_{0})$. If $\bar{e}\in\bar{\mathcal{L}}_{O_{M}(\mathcal{K})}$ and $\bar{f}\in\bar{\mathcal{L}}_{sep}$ are the images of $e$ and, resp., of $f$ under the extensions of scalars of the projection $\mathcal{L}\to\bar{\mathcal{L}}$ then $\tau\mapsto(-\bar{f})\circ\tau(\bar{f})$ determines the projection $\bar{\eta}^{0}:\mathcal{G}_{<p,M}\to G(\bar{\mathcal{L}})$. Clearly, $\bar{\eta}^{0}$ is induced by $\eta^{0}$.
As earlier, choose $r^{*}$, $N^{*}$, set $q=p^{N^{*}}$, $b^{*}=r^{*}(q-1)$, introduce the system of coefficients $\{t^{-\iota}\ |\ \iota\in\mathfrak{A}^{0}\}$ and consider the Lie algebra $\mathcal{L}^{{\dagger}}$. Then $\mathcal{L}^{{\dagger}}_{W(k)}$ has generators $D^{{\dagger}}_{\iota n}$, where $\iota\in\mathfrak{A}^{0}(p)$. By Proposition 2.4 all $\iota p^{-v_{p}(\iota)}$ are pairwise different and, therefore, we can define isomorphic embedding $\psi:\mathcal{L}^{{\dagger}}\to\mathcal{L}$ by the correspondences $D^{{\dagger}}_{\iota n}\mapsto D_{\iota^{*}n^{*}}$, where $\iota^{*}=\iota p^{-v_{p}(\iota)}$ and $n^{*}=n+v_{p}(\iota)$.
Let $\psi(\mathcal{L}^{{\dagger}})=\mathcal{L}^{o{\dagger}}$. If $\mathrm{ch}(\iota)=0$ then $\iota=aqp^{M-1}$, where $a\in[0,(p-1)v_{0})$, $\iota^{*}=a$ and $\psi(D^{{\dagger}}_{\iota n})=D_{a,n+M-1}$. If $\mathrm{ch}(\iota)\geqslant 1$ then $\iota p^{-v_{p}(\iota)}>(p-1)v_{0}$, cf. Prop. 2.4. In particular, for all such $\iota$, $D_{\iota n}\in\mathcal{L}(p)_{W(k)}$.
As earlier, define the filtration on $\mathcal{L}^{o{\dagger}}$ and set ¯$\bar{\mathcal{L}}^{o{\dagger}}=\mathcal{L}^{o{\dagger}}/\mathcal{L}^{o{\dagger}}(p)$. By Prop. 3.4, if $\mathrm{ch}(\iota)\geqslant 1$ then all $D_{\iota n}\in\bar{\mathcal{L}}(p)_{W(k)}$, and we have the induced projection $\bar{\mathcal{V}}^{o}:\bar{\mathcal{L}}^{o{\dagger}}\to\bar{\mathcal{L}}$.
@@
$e^{o{\dagger}}=\sum_{\iota\in\mathfrak{A}^{+}(p)}t^{-\iota}D_{\iota^{*}n^{*}}\,.$
Then the image of $\bar{e}^{o{\dagger}}$ in ¯$\bar{\mathcal{L}}_{O_{M}(\mathcal{K})}$ coincides with $\sigma^{M-1+N^{*}}(\bar{e})$. Therefore, we can choose $f^{{\dagger}}\in\bar{\mathcal{L}}^{{\dagger}}_{sep}$ such that its image in ¯$\bar{\mathcal{L}}_{sep}$ coincides with $\sigma^{M-1+N^{*}}\bar{f}$. This means that we have epimorphism $\mathcal{G}_{<p,M}\to G(\bar{\mathcal{L}}^{o{\dagger}})$ which gives $\bar{\eta}^{0}$ when composed with projection $G(\bar{\mathcal{V}^{o}}):G(\bar{\mathcal{L}}^{o{\dagger}})\to G(\bar{\mathcal{L}})$.
Now Theorem 3.1 states that a unipotent $\mathbb{Z}/p^{M}$-action determined by $t^{qp^{M-1}}\mapsto t^{qp^{M-1}}\widetilde{\exp}(-t^{b^{*}})^{p^{M-1}}$ induces via formalism of nilpotent Artin-Schreier theory a deformation $B^{o{\dagger}}:\mathrm{Ker}\,\bar{\mathcal{V}}^{o}\to\mathrm{Ker}\,\bar{\mathcal{V}}^{o}$. This deformation allows us to recover the “substantial” part ¯$\bar{\mathcal{L}}^{(v_{0})}$ of the ramification ideal $\mathcal{L}^{(v_{0})}$. It appears as $\bar{\mathcal{V}}^{o}(\mathrm{Ker}\,\bar{\mathcal{V}}^{o})$, i.e. as a deformation of the “trivial” part $\mathrm{Ker}\,\bar{\mathcal{V}}^{o}\subset\mathcal{L}(p)$ via the nilpotent operator $B^{o{\dagger}}$.
Such interpretation also indicates a chance of developing a theory which works for all $v_{0}$ (rather than just for only one single value $v_{0}$).
@@ 4.1. Agreements
We need to treat the Campbell-Hausdorff group law via exponentials in the appropriate enveloping algebras $\bar{\mathcal{A}}$ and $\bar{\mathcal{A}}^{{\dagger}}$ of $\bar{\mathcal{L}}$ and, resp., $\bar{\mathcal{L}}^{{\dagger}}$, by involving extension of scalars to $O^{\prime}_{M}(\mathcal{K}^{\prime})$. As a matter of fact, the definition of the action $\Omega$ requires working inside $\widetilde{\mathcal{N}}^{{\dagger}}$ but can be formally treated in $\bar{\mathcal{N}}^{{\dagger}}$, cf. Prop. 2.14. So, it makes sense to work mainly with the Lie algebra ¯$\bar{\mathcal{L}}^{{\dagger}}[0]\simeq\bar{\mathcal{L}}$ and its enveloping algebra $\bar{\mathcal{A}}^{{\dagger}}[0]$.
We are going also to use the advantage of special coefficients from $\mathfrak{A}^{0}$. So, we need just a part of the quotient $(\bar{\mathcal{A}}^{{\dagger}}/\bar{\mathcal{J}}^{{\dagger}p})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ generated by the elements of $\bar{\mathcal{N}}^{{\dagger}}[0]$. (Here $\bar{\mathcal{J}}^{{\dagger}}$ is the augmentation ideal of $\bar{\mathcal{A}}^{{\dagger}}$.) Even more, the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ is completely determined by the elements $\bar{\mathcal{V}}B^{{\dagger}}(D^{\,{\dagger}}_{\iota 0})$ with $\mathrm{ch}(\iota)=1$. Therefore, it would be reasonable to ignore all $\iota$ with $\mathrm{ch}(\iota)\geqslant 2$, i.e. to work with the part generated by $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}[0]$ consisting of the preimages of $\widetilde{\mathcal{N}}^{\,{\dagger}}[0,0]\oplus\widetilde{\mathcal{N}}^{{\dagger}}[0,1]$.
As a result we are going to perform calculations inside the $W_{M}(k)$-module $\bar{\mathcal{A}}^{\,{\dagger}}_{\leqslant 1}[0]$ generated by all $p^{A}t_{1}^{-\iota}D_{\bar{a}\bar{n}}$, where for $\bar{a}=(a_{1},\dots,a_{r})$, $\bar{n}=(n_{1},\dots,n_{r})$, $D_{\bar{a}\bar{n}}=D_{a_{1}n_{1}}\ldots D_{a_{r}n_{r}}$, $\iota\in\mathfrak{A}^{0}$, $\mathrm{ch}(\iota)\in\{0,1\}$ and it holds $-\iota\geqslant-qp^{M-1}|\bar{a}|+p^{M-1-A}b^{*}$ if $\mathrm{ch}(\iota)=1$ (recall that $U(1,A)=p^{-A}$). Since we do not need to follow the effects related to $\mathrm{ch}(\iota)\geqslant 2$, we can agree that $t_{1}^{-\iota}\cdot t_{1}^{-\iota^{\prime}}=0$ if $\mathrm{ch}(\iota)=\mathrm{ch}(\iota^{\prime})=1$. We do need to go to $\bar{\mathcal{A}}^{{\dagger}}_{\leqslant 1}[0]$ after applying exponential map to $\bar{\mathcal{N}}^{{\dagger}}_{\leqslant 1}$ but almost immediately we come back to $\bar{\mathcal{N}}^{{\dagger}}[0]$ and even to $\bar{\mathcal{N}}^{{\dagger}}[0,1]$.